OpenStudy (anonymous):

6x + 4x = 42 -3x + 3y = -6 #elimination, help? :)

OpenStudy (anonymous):

is that supposed to be 6x+4y?

OpenStudy (anonymous):

mulitply 2nd part by (3) then solve for y

OpenStudy (anonymous):

yesss matt

OpenStudy (anonymous):

6x+4y =42 2(-3x+3y = -6) 6x+4y=42 -6x+6y=-12 add both equations 10y=30 y=10 plug in 10 for y in either equation to solve for x